OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3. y - 6 = -3x - 3 y - 6 = -3(x + 1) y = -3x + 3 3x + y = 3

OpenStudy (solomonzelman):

plug in your point (-1,6) , which one works? It is B y - 6 = -3(x + 1) (6) - 6 = -3((-1) + 1) 6-6=-3(0) 0=0